OpenStudy (anonymous):

please answer S = p/(q + p(1 − q)) solve for q i don't know what to do.

jimthompson5910 (jim_thompson5910):

S = p/(q+p(1-q)) S(q+p(1-q)) = p S(q+p-pq) = p Sq+Sp-Spq = p Sq-Spq = p-Sp q(S-Sp) = p-Sp q = (p-Sp)/(S-Sp)

OpenStudy (anonymous):

wow thank you so much! very clear